OpenStudy (anonymous):

Solve each equation by completing the square. If necessary, round to the nearest hundredth.

OpenStudy (anonymous):

31. x^2=3x-1 33. x^2+10x=-8 35. x^2-14x+16=0

OpenStudy (anonymous):

Check out this link for an example on how to solove these types of problems. http://www.purplemath.com/modules/sqrquad.htm

OpenStudy (anonymous):

31) x=12(3+5√),12(3−5√) x≈2.61803398875,0.38196601125

OpenStudy (anonymous):

Thank you. What about the others?

OpenStudy (anonymous):

https://www.mathway.com go to this site it'll give you the answers.